OpenStudy (anonymous):

I am taking linear algebra. The question asks me to express in the form a+ib for the following 1) -1(i-1) 2) i^5 3) i^n 4) (1+i)^2 5) (i^3(1+i)^5)/(1+6)^6 6) (i(i^2)(i^3)..(i^n))/(i(i^3)(i^5)..(i^(2n+1))) Please any help would be greatly appreciated. Also, if possible please explain or write out the work so I know how to get the answer. Thank you so much

OpenStudy (anonymous):

1) -1(i-1) = 1 - i

OpenStudy (anonymous):

now how do i express it in the form a+ib, is it (1)+i(-1)?

OpenStudy (anonymous):

Recognize that a = constant, b is the imaginary part!

OpenStudy (anonymous):

a = 1, b = -1 => 1 - i

OpenStudy (anonymous):

okay, i wasn't sure if i was doing this properly

OpenStudy (anonymous):

For example \[ i^2 = -1\\ i^3= -i\\ i^4= 1\\ i^5= i = 0 +1 i\\ \]